The current stock price for a company is $39 per share, and there are 8 million shares outstanding. The beta for this firms stock is 1.3, the risk-free rate is 4.5, and the expected market risk premium is 5.9%. This firm also has 130,000 bonds outstanding, which pay interest semiannually. These bonds have a coupon interest rate of 9%, 28 years to maturity, a face value of $1,000, and a current price of 971.59. If the corporate tax rate is 38%, what is the Weighted Average Cost of Capital (WACC) for this firm? (Answer to the nearest hundredth of a percent, but do not use a percent sign).
Answer #1

Value of debt = 130000*971.59 = 126306700

Value of equity = 39*8000000 = 312000000

Total value = 126306700 + 312000000 = 438306700

Cost of debt:

PV = 971.59

FV = 1000

N = 28*2 = 56

PMT = 5.9%/2*1000 = 29.5

Using fincancial calculator, I/Y = 3.057%

Annual YTM = 3.057%*2 = 6.11%

Cost of equity:

Cosst of equity = Risk free rate + beta*Risk premium = 4.5%+1.3*5.9% = 12.17%

WACC = rD (1- Tc )*( D / V )+ rE *( E / V )

Where...

rD = The required return of the firm's Debt financing
(1-Tc) = The Tax adjustment for interest expense
(D/V) = (Debt/Total Value)
rE= the firm's cost of equity
(E/V) = (Equity/Total Value)

WACC = 126306700/438306700 * 6.11%*(1-0.38) + 312000000/438306700*12.17% = 9.75%

WACC = 9.75% = 0.0975


answered by: ANURANJAN SARSAM